find the slope of a line perpendicular to the line that passes through(3,9) and (7,15)
Answer by mananth(16946)

x1 y1 x2 y2
3 * 9 7 15

slope m = (y2-y1)/(x2-x1)
( 15 - 9 )/( 7 - 3 )
( 6 / 4 )
m= 3/ 2
The slope of perpendicular line = -2/3 because m1*m2 = -1